### v1

@@ 1 Introduction
Marek Jarnicki was a great expert on invariant distances and metrics in complex analysis and geometry; his book with Peter Pflug [JarPfl2013] is a definitive reference on the subject. Since the 1920’s, invariant distances have often been instrumental in the study of holomorphic dynamics on hyperbolic Riemann surfaces and hyperbolic manifolds; a good example is Wolff’s use [Wol1926] of the result that has later become known as the Wolff lemma to simplify the proof of the Wolff-Denjoy theorem (see Theorem 1.1) describing the dynamics of holomorphic self-maps of the unit disk $\mathbb{D}$ of the complex plane.
Maurice Heins (1915–2015) was a master in the study of holomorphic dynamics on hyperbolic Riemann surfaces; for instance, he has been able [He1941] to extend the Wolff-Denjoy theorem to multiply connected domains in the plane (for more on this beautiful theory and its history see, e.g., [Ab2022] and references therein).
He often used invariant distances in his work, as confirmed by a recent unexpected (at least by me) discovery. While Ian Short and myself were working on the paper [AbSh2025] on iterated function systems (where the object of study is the dynamics of a sequence of functions obtained by composing different self-maps), we came across (actually, Ian did) an unpublished manuscript by Heins [He1991], where with a clever use of the Poincaré invariant distance he proved a number of beautiful results on holomorphic dynamics in $\mathbb{D}$. We incorporated, extended and generalized some of these results in [AbSh2025], but a few of them were left out, because not relevant to our setting. In particular, we left out an elegant solution to a question on parabolic dynamics posed by Valiron in [Val1954]. This question has already been answered by Pommerenke in [Pomm1979], but Heins’ approach is surprisingly simpler and based on ideas having other applications too (see, for instance, [AbSh2025] and Section 2 of this paper).
I believe that Heins’ ideas deserve to be better known; being based on invariant distances, they fit well in a publication dedicated to Marek. So this paper is a report on Heins’ approach to parabolic dynamics as described in [He1991]. I have updated the presentation, in order to make it coherent with contemporary research on this topic, and streamlined some proofs; but the main ideas are his.
@@
If $f\in\operatorname{\textnormal{Hol}}(\mathbb{D},\mathbb{D})$ has a fixed point $z_{0}\in\mathbb{D}$, an easy application of the Schwarz-Pick lemma shows that either $f$ is an automorphism (and then its dynamics is the same as the dynamics of an Euclidean rotation) or the sequence of iterates converges to $z_{0}$. Furthermore, thanks to the work of Kœnigs [Koe] in 1883 and Böttcher [Bot] in 1904, we know exactly how the orbits converge to the fixed point.
@@
The point $\tau_{f}$ is the Wolff point of the map $f\in\operatorname{\textnormal{Hol}}(\mathbb{D},\mathbb{D})\setminus\operatorname{\textnormal{Aut}}(\mathbb{D})$. Using the position of the Wolff point and the value of $f^{\prime}(\tau_{f})$ we can introduce the following classification: a holomorphic map $f\in\operatorname{\textnormal{Hol}}(\mathbb{D},\mathbb{D})$ is
  1. (i)
@@
The dynamics of hyperbolic maps is more complicated than the dynamics of elliptic maps but still well-behaved: Wolff [Wol1929] in 1929 and Valiron [Val1931] in 1931 proved that for any $z_{0}\in\mathbb{D}$ the orbit $\{f^{n}(z_{0})\}$ converges to the Wolff point non-tangentially and with a precise slope. More precisely, Valiron [Val1931] (see also [Val1954]) proved that if $z_{0}\in\mathbb{D}$ is fixed then there exists a non-constant holomorphic map $\psi$ such that
$\lim_{n\to+\infty}\frac{f^{n}(z)-\tau_{f}}{f^{n}(z_{0})-\tau_{f}}=\psi(z)$ (1.1)
@@
$\lim_{n\to+\infty}\frac{f^{n}(z)-\tau_{f}}{|f^{n}(z)-\tau_{f}|}=\tau_{f}e^{i\theta_{z}}\;.$ (1.2)
In the parabolic case, Wolff and Valiron were able to obtain partial results only. So in [Val1954, p. 148] Valiron asked whether the limit of the left-hand side in (1.1) does exist in the parabolic case too. Moreover, he wondered whether there are conditions ensuring that parabolic orbits converge to the Wolff point along a definite slope, like in (1.2). (Actually, Valiron worked in the right half-plane; the previous statements are the translation of Valiron’s results and questions to $\mathbb{D}$, which is biholomorphic to a half-plane in $\mathbb{C}$ via the Cayley transform; see Section 3.)
As mentioned before, Pommerenke [Pomm1979] answered these questions in 1979. However, his answers are a byproduct of quite a delicate argument whose aim is to find a solution to the Abel equation $\psi\circ f=\psi+b$ or, in modern terminology, to find models for parabolic maps. Here, a model is an automorphism of $\mathbb{D}$ or $\mathbb{C}$ which is semiconjugate in an appropriate sense to the original map $f$ and thus can be used to study its dynamics. The theory of models is very powerful but definitely not easy; see [Ab2022, Chapters 3 and 4] for an introduction.
It is worthwhile to remark immediately that Wolff and Valiron already knew that (1.2) does not hold for all parabolic maps. More precisely, Pommerenke clarified that the class of parabolic maps should be subdivided in two subclasses, having quite a different dynamical behaviour. In modern terminology, the subdivision is expressed in terms of the hyperbolic step, that, in turn, is defined by using the Poincaré distance $\omega$ on $\mathbb{D}$. Take $f\in\operatorname{\textnormal{Hol}}(\mathbb{D},\mathbb{D}))$. By the Schwarz-Pick theorem, for any $z\in\mathbb{D}$ the sequence $\bigl{\{}\omega\bigl{(}f^{n}(z),f^{n+1}(z)\bigr{)}\bigr{\}}$ is non-increasing and, thus, it is converging. The hyperbolic step $s^{f}\colon\mathbb{D}\to\mathbb{R}^{+}$ of $f$ is then defined by
@@
It turns out that the dynamics of parabolic maps with positive hyperbolic step (that is, such that $s^{f}>0$ everywhere) is quite different from the dynamics of parabolic maps with zero hyperbolic step (that is, such that $s^{f}\equiv 0$). In particular, (1.2) holds for parabolic maps with positive hyperbolic step but there are examples of parabolic maps with zero parabolic step where (1.2) does not hold.
In principle, there might be a third class of parabolic maps, consisting of maps for which the hyperbolic step vanishes somewhere but it is not identically zero. However, it turns out that this is not the case: $s^{f}$ vanishes at one point if and only if it is identically zero. This result, which is instrumental in proving (1.2) for parabolic maps with positive hyperbolic step, has been obtained by using the theory of models; however, in [Ab2022, p. 248] I wondered whether it was possible to give a proof independent of the theory of models.
Heins’ answers to these questions is based on the apparently unrelated idea of left straightening of the sequence of iterates $\{f^{n}\}$ of a map $f$ (here I am using the terminology introduced, in a more general setting, in [AbSh2025]). A left straightening of $\{f^{n}\}$ is a holomorphic self-map $h\in\operatorname{\textnormal{Hol}}(\mathbb{D},\mathbb{D})$ obtained as limit of a sequence of the form $\{\gamma_{n}^{-1}\circ f^{n}\}$, with $\gamma_{n}\in\operatorname{\textnormal{Aut}}(\mathbb{D})$ for all $n\in\mathbb{N}$.
In Section 2 we shall prove that a left straightening always exists and is unique up to left composition by an automorphism (Theorem 2.1). Since the automorphisms of $\mathbb{D}$ are isometries for the Poincaré distance, a left straightening $h$ of $\{f^{n}\}$ can be used to express the hyperbolic step: indeed, we have $s^{f}(z)=\omega\bigl{(}h(z),h\bigl{(}f(z)\bigr{)}$ for all $z\in\mathbb{D}$ (Corollary 2.2). So $s^{f}(z_{0})=0$ if and only if $h\bigl{(}f(z_{0})\bigr{)}=h(z_{0})$. Having this, a clever application of Rouché’s theorem allows to prove that $s^{f}$ vanishes at one point if and only if it is identically zero (Theorem 2.3).
Finally, in Section 3 we shall show how to use the hyperbolic step and Theorem 2.3 to push the original Valiron’s argument just that little bit further needed to answer his questions: when $f$ is parabolic, then the ratio in (1.1) has limit 1 (Corollary 3.6) and when $p$ is parabolic with positive hyperbolic step then the ratio in (1.2) has the same limit  $\pm i\tau_{f}$ for all $z\in\mathbb{D}$ (Corollary 3.10). The final argument is deceptively simple; but this is just a consequence of the elegance and strength of Heins’ ideas.
@@ 2 Left straightening and hyperbolic step
Let us fix a few notations and recall a few well-know results; the details can be found in [Ab2022] and references therein.
@@
$\omega\bigl{(}f(z),f(w)\bigr{)}\leqslant\omega(z,w)$
@@
In this section we shall give a simple proof of this fact, by using the idea of left straightening of a sequence of holomorphic self-maps, implicitly introduced by Heins in the unpublished manuscript [He1991] for the sequence of iterates of a single map and then generalized by Short and the author [AbSh2025] to arbitrary iterated function systems of $\mathbb{D}$. For the sake of completeness, we report here a proof of the existence of a left straightening of a sequence of iterates (for the general case see [AbSh2025, Theorem A]). The original proof by Heins was based on Harnack convergence theorem; here we instead use a normal families argument, which is more in line with the rest of the paper.
Theorem 2.1.
@@
Since $\omega\bigl{(}f^{n}(z),f^{n}(0)\bigr{)}\leqslant\omega\bigl{(}f^{n-1}(z),f^{n-1}(0))$, it follows that $\omega\bigl{(}H_{n}(z),0\bigr{)}\leqslant\omega\bigl{(}H_{n-1}(z),0\bigr{)}$; therefore $\{|H_{n}|\}$ is a non-increasing sequence. If $|H_{n}|\to 0$ (pointwise and hence, by Vitali’s theorem, uniformly on compact subsets), then we can take $h\equiv 0$ and we are done. Otherwise, there exists $w_{0}\in\mathbb{D}$ for which $\{|H_{n}(w_{0})|\}$ converges to a positive constant. For any $n\in\mathbb{N}$, let $\theta_{n}$ be an argument of $H_{n}(w_{0})$. By pre-composing $\gamma_{n}$ with the rotation $z\mapsto e^{i\theta_{n}}z$, we can assume that $\{H_{n}(w_{0})\}$ is a non-increasing sequence of positive numbers converging to a positive number $\rho_{0}\in(0,1)$.
@@
$H_{n_{i}}=K_{i}\circ H_{m_{i}}\;.$ (2.1)
Note that $K_{i}(z_{0})=0$; so $\{K_{i}\}$ is relatively compact in $\operatorname{\textnormal{Hol}}(\mathbb{D},\mathbb{D})$. Consequently, there is a subsequence of $\{K_{i}\}$ converging to $\psi\in\operatorname{\textnormal{Hol}}(\mathbb{D},\mathbb{D})$; from (2.1) we infer that $\psi\circ h=k$. Notice that
@@
$\omega\bigl{(}\phi_{n}\bigl{(}k(z_{0})\bigr{)},h(z_{0})\bigr{)}$ $\leqslant\omega\bigl{(}\phi_{n}\bigl{(}k(z_{0})\bigr{)},\phi_{n}\bigl{(}\delta_{n}^{-1}\bigl{(}f^{n}(z_{0})\bigr{)}\bigr{)}\bigr{)}+\omega\bigl{(}\phi_{n}\bigl{(}\delta_{n}^{-1}\bigl{(}f^{n}(z_{0})\bigr{)}\bigr{)},h(z_{0})\bigr{)}$
@@
In [AbSh2025] we used the left straightening to prove several results on iterated function systems. Here, the main point is that we can use any left straightening to compute the hyperbolic step:
Corollary 2.2.
@@
$\lim_{n\to+\infty}\omega\bigl{(}f^{n}(z),f^{n}(w)\bigr{)}=\omega\bigl{(}h(z),h(w)\bigr{)}\;.$ (2.2)
@@
for any $z$, $w\in\mathbb{D}$ and (2.2) is proved. The final assertion follows immediately by taking $w=f(z)$. ∎
@@
That (ii) implies (iii) follows immediately from (2.2). Taking $w=f(z)$ we see that (iii) implies (iv). Moreover, (iv) trivially implies (i); so we are left with proving that (i) implies (ii).
@@
$\leqslant\omega\bigl{(}f^{n-1}(z_{0}),f^{n-1}(z)\bigr{)}=\omega\bigl{(}0,H_{n-1}(z)\bigr{)}\;.$
So $\{|H_{n}|\}$ converges non-increasingly to $|h|$; in particular, $|H_{n}|\geqslant|h|$ for all $n\in\mathbb{N}$.
Next, by Corollary 2.2, $s^{f}(z_{0})=0$ if and only if $\omega\bigl{(}h(z_{0}),h\bigl{(}f(z_{0})\bigr{)}\bigr{)}=0$; since $h(z_{0})=0$, we must have $h\bigl{(}f(z_{0})\bigr{)}=0$. Notice that $f$ has no fixed points; so $f(z_{0})$ is a distinct zero of $h$. We claim that then $h\equiv 0$.
Assume, by contradiction, that $h$ is not identically zero. Then $f(z_{0})$ is an isolated zero of $h$; choose a small $r>0$ so that, if $B=B\bigl{(}f(z_{0}),r\bigr{)}\subset\mathbb{D}$ is the Euclidean ball of center $f(z_{0})$ and radius $r$, then $f(z_{0})$ is the unique zero of $h$ in $\overline{B}$. Let $\varepsilon=\inf\limits_{\zeta\in\partial B}|h(\zeta)|>0$. Since $H_{n}\to h$ uniformly on compact subsets, we have $\sup\limits_{\zeta\in\partial B}|H_{n}(\zeta)-h(\zeta)|<\varepsilon$ for $n$ large enough. By Rouché theorem, $H_{n}$ and $h$ must eventually have the same number of zeroes in $B$, counted with multiplicities. But $h$ in $B$ vanishes only in $f(z_{0})$; since $|H_{n}|\geqslant|h|$, the same must happen for $H_{n}$ too as soon as $n$ is large enough.

### v2

@@ Abstract
keywords:
holomorphic self-map, holomorphic dynamics, parabolic maps, hyperbolic step, left straightening.
pacs:
MSC Classification]30D05 37F99
@@
[…]
Acknowledgements The author was partially supported by the French-Italian University and by Campus France through the Galileo program, under the project From rational to transcendental: complex dynamics and parameter spaces, as well as by Istituto Nazionale di Alta Matematica (INdAM).
@@ 1 Introduction
Marek Jarnicki was a great expert on invariant distances and metrics in complex analysis and geometry; his book with Peter Pflug [1] is a definitive reference on the subject. Since the 1920’s, invariant distances have often been instrumental in the study of holomorphic dynamics on hyperbolic Riemann surfaces and hyperbolic manifolds; a good example is Wolff’s use [2] of the result that has later become known as the Wolff lemma to simplify the proof of the Wolff-Denjoy theorem (see Theorem 1) describing the dynamics of holomorphic self-maps of the unit disk $\mathbb{D}$ of the complex plane.
Maurice Heins (1915–2015) was a master in the study of holomorphic dynamics on hyperbolic Riemann surfaces; for instance, he has been able [3] to extend the Wolff-Denjoy theorem to multiply connected domains in the plane (for more on this beautiful theory and its history see, e.g., [4] and references therein).
He often used invariant distances in his work, as confirmed by a recent unexpected (at least by me) discovery. While Ian Short and myself were working on the paper [5] on iterated function systems (where the object of study is the dynamics of a sequence of functions obtained by composing different self-maps), we came across (actually, Ian did) an unpublished manuscript by Heins [6], where with a clever use of the Poincaré invariant distance he proved a number of beautiful results on holomorphic dynamics in $\mathbb{D}$. We incorporated, extended and generalized some of these results in [5], but a few of them were left out, because they were not relevant to our setting. In particular, we left out an elegant solution to a question on parabolic dynamics posed by Valiron in [7]. This question has already been answered by Pommerenke in [8], but Heins’ approach is surprisingly simpler and based on ideas having other applications too (see, for instance, [5] and Section 2 of this paper).
I believe that Heins’ ideas deserve to be better known; being based on invariant distances, they fit well in a publication dedicated to Marek. So this paper is a report on Heins’ approach to parabolic dynamics as described in [6]. I have updated the presentation, in order to make it coherent with contemporary research on this topic, and streamlined some proofs; but the main ideas are his.
@@
If $f\in\operatorname{\textnormal{Hol}}(\mathbb{D},\mathbb{D})$ has a fixed point $z_{0}\in\mathbb{D}$, an easy application of the Schwarz-Pick lemma shows that either $f$ is an automorphism (and then its dynamics is the same as the dynamics of an Euclidean rotation) or the sequence of iterates converges to $z_{0}$. Furthermore, thanks to the work of Kœnigs [9] in 1883 and Böttcher [10] in 1904, we know exactly how the orbits converge to the fixed point.
@@
The point $\tau_{f}$ is the Wolff point of the map $f$. Using the position of the Wolff point and the value of $f^{\prime}(\tau_{f})$ we can introduce the following classification: a holomorphic map $f\in\operatorname{\textnormal{Hol}}(\mathbb{D},\mathbb{D})$ is
  1. 1.
@@
The dynamics of hyperbolic maps is more complicated than the dynamics of elliptic maps but still well-behaved: Wolff [11] in 1929 and Valiron [12] in 1931 proved that for any $z_{0}\in\mathbb{D}$ the orbit $\{f^{n}(z_{0})\}$ converges to the Wolff point non-tangentially and with a precise slope. More precisely, Valiron [12] (see also [7]) proved that if $z_{0}\in\mathbb{D}$ is fixed then there exists a non-constant holomorphic map $\psi$ such that
$\lim_{n\to+\infty}\frac{f^{n}(z)-\tau_{f}}{f^{n}(z_{0})-\tau_{f}}=\psi(z)$ (1)
@@
$\lim_{n\to+\infty}\frac{f^{n}(z)-\tau_{f}}{|f^{n}(z)-\tau_{f}|}=\tau_{f}e^{i\theta_{z}}\;.$ (2)
In the parabolic case, Wolff and Valiron were able to obtain partial results only. So in [7, p. 148] Valiron asked whether the limit of the left-hand side in (1) does exist in the parabolic case too. Moreover, he wondered whether there are conditions ensuring that parabolic orbits converge to the Wolff point along a definite slope, like in (2). (Actually, Valiron worked in the right half-plane; the previous statements are the translation of Valiron’s results and questions to $\mathbb{D}$, which is biholomorphic to a half-plane in $\mathbb{C}$ via the Cayley transform; see Section 3.)
As mentioned before, Pommerenke [8] answered these questions in 1979. However, his answers are a byproduct of quite a delicate argument whose aim is to find a solution to the Abel equation $\psi\circ f=\psi+b$ or, in modern terminology, to find models for parabolic maps. Here, a model is an automorphism of $\mathbb{D}$ or $\mathbb{C}$ which is semiconjugate in an appropriate sense to the original map $f$ and thus can be used to study its dynamics. The theory of models is very powerful but definitely not easy; see [4, Chapters 3 and 4] for an introduction.
It is worthwhile to remark immediately that Wolff and Valiron already knew that (2) does not hold for all parabolic maps. More precisely, Pommerenke clarified that the class of parabolic maps should be subdivided in two subclasses, having quite a different dynamical behaviour. In modern terminology, the subdivision is expressed in terms of the hyperbolic step, that, in turn, is defined by using the Poincaré distance $\omega$ on $\mathbb{D}$. Take $f\in\operatorname{\textnormal{Hol}}(\mathbb{D},\mathbb{D})$. By the Schwarz-Pick theorem, for any $z\in\mathbb{D}$ the sequence $\bigl\{\omega\bigl(f^{n}(z),f^{n+1}(z)\bigr)\bigr\}$ is non-increasing and, thus, it is converging. The hyperbolic step $s^{f}\colon\mathbb{D}\to\mathbb{R}^{+}$ of $f$ is then defined by
@@
It turns out that the dynamics of parabolic maps with positive hyperbolic step (that is, such that $s^{f}>0$ everywhere) is quite different from the dynamics of parabolic maps with zero hyperbolic step (that is, such that $s^{f}\equiv 0$). In particular, (2) holds for parabolic maps with positive hyperbolic step but there are examples of parabolic maps with zero parabolic step where (2) does not hold.
In principle, there might be a third class of parabolic maps, consisting of maps for which the hyperbolic step vanishes somewhere but it is not identically zero. However, it turns out that this is not the case: $s^{f}$ vanishes at one point if and only if it is identically zero. This result, which is instrumental in proving (2) for parabolic maps with positive hyperbolic step, has been obtained by using the theory of models; however, in [4, p. 248] I wondered whether it was possible to give a proof independent of the theory of models.
Heins’ answers to these questions is based on the apparently unrelated idea of left straightening of the sequence of iterates $\{f^{n}\}$ of a map $f$ (here I am using the terminology introduced, in a more general setting, in [5]). A left straightening of $\{f^{n}\}$ is a holomorphic self-map $h\in\operatorname{\textnormal{Hol}}(\mathbb{D},\mathbb{D})$ obtained as limit of a sequence of the form $\{\gamma_{n}^{-1}\circ f^{n}\}$, with $\gamma_{n}\in\operatorname{\textnormal{Aut}}(\mathbb{D})$ for all $n\in\mathbb{N}$.
In Section 2 we shall prove that a left straightening always exists and is unique up to left composition by an automorphism (Theorem 2). Since the automorphisms of $\mathbb{D}$ are isometries for the Poincaré distance, a left straightening $h$ of $\{f^{n}\}$ can be used to express the hyperbolic step: indeed, we have $s^{f}(z)=\omega\bigl(h(z),h\bigl(f(z)\bigr)$ for all $z\in\mathbb{D}$ (Corollary 3). So $s^{f}(z_{0})=0$ if and only if $h\bigl(f(z_{0})\bigr)=h(z_{0})$. Having this, a clever application of Rouché’s theorem allows to prove that $s^{f}$ vanishes at one point if and only if it is identically zero (Theorem 4).
Finally, in Section 3 we shall show how to use the hyperbolic step and Theorem 4 to push the original Valiron’s argument just that little bit further needed to answer his questions: when $f$ is parabolic, then the ratio in (1) has limit 1 (Corollary 8) and when $p$ is parabolic with positive hyperbolic step then the ratio in (2) has the same limit  $\pm i\tau_{f}$ for all $z\in\mathbb{D}$ (Corollary 12). The final argument is deceptively simple; but this is just a consequence of the elegance and strength of Heins’ ideas.
@@ 2 Left straightening and hyperbolic step
Let us fix a few notations and recall a few well-know results; the details can be found in [4] and references therein.
@@
$\omega\bigl(f(z),f(w)\bigr)\leq\omega(z,w)$
@@
In this section we shall give a simple proof of this fact, by using the idea of left straightening of a sequence of holomorphic self-maps, implicitly introduced by Heins in the unpublished manuscript [6] for the sequence of iterates of a single map and then generalized by Short and the author [5] to arbitrary iterated function systems of $\mathbb{D}$. For the sake of completeness, we report here a proof of the existence of a left straightening of a sequence of iterates (for the general case see [5, Theorem A]). The original proof by Heins was based on the Harnack convergence theorem; here we instead use a normal families argument, which is more in line with the rest of the paper.
Theorem 2.
@@
Since $\omega\bigl(f^{n}(z),f^{n}(0)\bigr)\leq\omega\bigl(f^{n-1}(z),f^{n-1}(0))$, it follows that $\omega\bigl(H_{n}(z),0\bigr)\leq\omega\bigl(H_{n-1}(z),0\bigr)$; therefore $\{|H_{n}|\}$ is a non-increasing sequence. If $|H_{n}|\to 0$ (pointwise and hence, by Vitali’s theorem, uniformly on compact subsets), then we can take $h\equiv 0$ and we are done. Otherwise, there exists $w_{0}\in\mathbb{D}$ for which $\{|H_{n}(w_{0})|\}$ converges to a positive constant. For any $n\in\mathbb{N}$, let $\theta_{n}$ be an argument of $H_{n}(w_{0})$. By pre-composing $\gamma_{n}$ with the rotation $z\mapsto e^{i\theta_{n}}z$, we can assume that $\{H_{n}(w_{0})\}$ is a non-increasing sequence of positive numbers converging to a positive number $\rho_{0}\in(0,1)$.
@@
$H_{n_{i}}=K_{i}\circ H_{m_{i}}\;.$ (3)
Note that $K_{i}(z_{0})=0$; so $\{K_{i}\}$ is relatively compact in $\operatorname{\textnormal{Hol}}(\mathbb{D},\mathbb{D})$. Consequently, there is a subsequence of $\{K_{i}\}$ converging to $\psi\in\operatorname{\textnormal{Hol}}(\mathbb{D},\mathbb{D})$; from (3) we infer that $\psi\circ h=k$. Notice that
@@
$\omega\bigl(\phi_{n}\bigl(k(z_{0})\bigr),h(z_{0})\bigr)$ $\leq\omega\bigl(\phi_{n}\bigl(k(z_{0})\bigr),\phi_{n}\bigl(\delta_{n}^{-1}\bigl(f^{n}(z_{0})\bigr)\bigr)\bigr)+\omega\bigl(\phi_{n}\bigl(\delta_{n}^{-1}\bigl(f^{n}(z_{0})\bigr)\bigr),h(z_{0})\bigr)$
@@
In [5] we used the left straightening to prove several results on iterated function systems. Here, the main point is that we can use any left straightening to compute the hyperbolic step:
Corollary 3.
@@
$\lim_{n\to+\infty}\omega\bigl(f^{n}(z),f^{n}(w)\bigr)=\omega\bigl(h(z),h(w)\bigr)\;.$ (4)
@@
for any $z$, $w\in\mathbb{D}$ and (4) is proved. The final assertion follows immediately by taking $w=f(z)$. ∎
@@
That (ii) implies (iii) follows immediately from (4). Taking $w=f(z)$ we see that (iii) implies (iv). Moreover, (iv) trivially implies (i); so we are left with proving that (i) implies (ii).
@@
$\leq\omega\bigl(f^{n-1}(z_{0}),f^{n-1}(z)\bigr)=\omega\bigl(0,H_{n-1}(z)\bigr)\;.$
So $\{|H_{n}|\}$ converges non-increasingly to $|h|$; in particular, $|H_{n}|\geq|h|$ for all $n\in\mathbb{N}$.
Next, by Corollary 3, $s^{f}(z_{0})=0$ if and only if $\omega\bigl(h(z_{0}),h\bigl(f(z_{0})\bigr)\bigr)=0$; since $h(z_{0})=0$, we must have $h\bigl(f(z_{0})\bigr)=0$. Notice that $f$ has no fixed points; so $f(z_{0})$ is a distinct zero of $h$. We claim that then $h\equiv 0$.
Assume, by contradiction, that $h$ is not identically zero. Then $f(z_{0})$ is an isolated zero of $h$; choose a small $r>0$ so that, if $B=B\bigl(f(z_{0}),r\bigr)\subset\mathbb{D}$ is the Euclidean ball of center $f(z_{0})$ and radius $r$, then $f(z_{0})$ is the unique zero of $h$ in $\overline{B}$. Let $\varepsilon=\inf\limits_{\zeta\in\partial B}|h(\zeta)|>0$. Since $H_{n}\to h$ uniformly on compact subsets, we have $\sup\limits_{\zeta\in\partial B}|H_{n}(\zeta)-h(\zeta)|<\varepsilon$ for $n$ large enough. By Rouché theorem, $H_{n}$ and $h$ must eventually have the same number of zeroes in $B$, counted with multiplicities. But $h$ in $B$ vanishes only in $f(z_{0})$; since $|H_{n}|\geq|h|$, the same must happen for $H_{n}$ too as soon as $n$ is large enough.